OpenStudy (anonymous):

How do you find Acceleration when they give you Nutens and Grams... (20N, 5g)

OpenStudy (anonymous):

Newton's 2nd Law states: F = ma You have been give the force 'F': F = 20 N You have been given the mass 'm': m = 5 g = 0.005 kg To find acceleration 'a', just divide: a = F / m = 20 / 0.005 = 4000 m/s^2

OpenStudy (mrnood):

it's NEWTONS Named after one of the most famous and innovative physicist of all times. Look up "Sir Isaac Newton"